OpenStudy (anonymous):

How many solutions does the equation −3y + 3y + 4 = 4 have?

OpenStudy (welshfella):

what do you think?

OpenStudy (welshfella):

try plugging in y = 1,2 and 3 and see what you get

OpenStudy (anonymous):

infinitly many

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

right

OpenStudy (anonymous):

The work of a student to solve the equation 3(3y − 8) = 10 + 3y is shown below: Step 1: 3(3y − 8) = 10 + 3y Step 2: 6y − 5 = 10 + 3y Step 3: 3y − 5 = 10 Step 4: 3y = 15 Step 5: y = 5 In which step did the student first make an error and what is the correct step?

OpenStudy (anonymous):

can you help me with one more?

OpenStudy (welshfella):

any real number can be substituted for y and the equation will still hold